### v1

@@ 4.3 Collision Cross Sections
As an illustrative example only, in our result for physical collisions among all LEO RSOs, we assume that the collision cross section for satellite-satellite, rocket body-satellite, rocket body-rocket body, debris-satellite, debris-rocket body, debris-debris be 300, 300, 300, 79, 79, 0.03 $\rm m^{2}$, respectively. Again, this is only intended to be approximate, yet meaningful. These values are also used for determining the current status of the CRASH Clock, but will need to be reevaluated as conditions and information change.
@@ 6 Code Availability
7 Acknowledgments
The authors would like to thank Christopher Chyba and Ryne Beeson for helpful discussions regarding our simulation code and the relative velocity calculation.
8 Author Contributions
All authors were fully involved in all aspects of this research. ST wrote the initial draft.
FUNDING Natural Sciences and Engineering Research Council of Canada DH-2022-00477 (AB, SH) and RGPIN-2020-04111 (SL)
9 Competing interests
Authors have no competing interests to declare.
10 Supplementary information
There is no supplementary material for this work.

### v2

@@ 4.3 Collision Cross Sections
As an illustrative reference example only, in our result for possible collisions among LEO RSOs we assume that the collision cross section for satellite-satellite, rocket body-satellite, rocket body-rocket body, debris-satellite, debris-rocket body, debris-debris be 300, 300, 300, 79, 79, 0.03 $\rm m^{2}$, respectively. These collisional areas, which are not the same as the individual cross section of each object, correspond to a close approach distance of 10 m between two satellites, satellite-rocket body, or two rocket bodies; 5 m between a satellite or rocket body and debris, and 10 cm between two pieces of debris. Again, this is only intended to be approximate, yet meaningful. These values are also used for determining the CRASH Clock in the main text, but this will need to be reevaluated as conditions and information change. In the article, we referred to this clock as the 10m-5m-10cm Clock. The motivation behind choosing these values is that they represent distances at which a collision could occur for many objects, depending on their orientation. These are thus close approach distances of serious concern – not guaranteed collisions.
Moreover, using the average area of a satellite could underestimate the collision potential. Take the Starlink V2 mini as an example, which has approximate dimensions of 0.3 m $\times$ 4.1 m $\times$ 29 m. The average surface area is approximately $43\rm~m^{2}$. If the collision cross section between two objects is defined as $A_{\rm col}=(A_{1}^{1/2}+A_{2}^{1/2})^{2}$, as is commonly done, the estimated collision cross section would be $61\rm m^{2}$. As discussed in the main text, two randomly oriented thin rods of full span $L$ have a typical collision distance (between rod centres) of approximately $0.3L$. Thus, for the Starlink V2 minis, assuming random orientations, the effective collision cross section approaches the $300\rm~m^{2}$ value we use – the average area of the satellite vastly underestimates the collision potential in this case.
Nevertheless, we are mindful that not all objects will have such geometries, and as such, it is important to explore different Clocks. Based on the ESA environmental report [ESA2025] of average megaconstellation satellite surface areas, a rough estimate of collisional encounter distances is 4.8 m (i.e., a collision cross section of about $72\rm~m^{2}$). Thus, if we take close approaches of 4.8 m, 2.4 m, and 10 cm (in the same way as above) to lead to collisions, the resulting 4.8m-2.4m-10cm Clock is 23 d for 2025 and 701 d for 2018.
In short, the Clock value is dependent on assumptions of the minimum tolerable close approaches, but the trends highlighted by various Clocks are ultimately the same.